OpenStudy (anonymous):

Find any relative extrema of the function. (Round your answers to three decimal places.) f(x) = arcsec(x) − 3x relative maximum (x, y) = ( ) relative minimum (x, y) = ( )

OpenStudy (anonymous):

can someone please help me

OpenStudy (dumbcow):

Differentiate the function and set equal to 0 hold on i have to look up the derivative of arcsec

OpenStudy (anonymous):

oki please thanx help me:(

OpenStudy (dumbcow):

\[d/dx arcsec(x) = \frac{1}{x \sqrt{x^{2}-1}}\]

OpenStudy (anonymous):

oki so were setting that to zero..umm how would u solve it from there

OpenStudy (dumbcow):

\[\frac{1}{x \sqrt{x^{2}-1}} - 3 = 0\]

OpenStudy (dumbcow):

\[\rightarrow x \sqrt{x^{2}-1} = \frac{1}{3}\]

OpenStudy (anonymous):

oki and then om gi have 5 mins left:(

OpenStudy (dumbcow):

this is timed? is this a test?

OpenStudy (anonymous):

yeh

OpenStudy (dumbcow):

haha good luck, this is not the best way to take a test, people aren't going to just give you quick answers, ....plus its kinda cheating

OpenStudy (anonymous):

i know but its just this oen problem i have trouble on

OpenStudy (dumbcow):

im sure well i gave you some help, all you have to do now is solve for x

OpenStudy (dumbcow):

ill give you next step square both sides